OpenStudy (anonymous):

10y-25=5 a.y=-3 b.y=-2 c.y=2 d.y=3

MsBrains (ms-brains):

D. :)

MsBrains (ms-brains):

y=3

OpenStudy (anonymous):

First,add 25 on both sides Then,divide both sides by 10

OpenStudy (like1234):

Let's solve your equation step-by-step. 10y−25=5 Step 1: Add 25 to both sides. 10y−25+25=5+25 10y=30 Step 2: Divide both sides by 10. 10y 10 = 30 10 y=3 Answer: y=3

OpenStudy (anonymous):

yep :)

OpenStudy (anonymous):

thank you